# Type 1 Diabetes Management using GLIMMER: Glucose Level Indicator Model with Modified Error Rate

### v1

@@ III-B Custom Loss Function
III-C Error Weights
We classify the CGM values according to the threshold values discussed previously to identify the respective regions. Our next step is to finalize the custom loss function by determining the optimal weight parameters for each region, with a primary focus on dysglycemia areas. For this reason and to simplify the optimization problem, we set to 1, reducing our optimization task to finding the optimal values for and .
To obtain optimal values of and , we propose to use Genetic Algorithms (GA) [45]. We chose this technique for several reasons. First, GAs are effective for optimization problems where the fitness function involves running another algorithm, particularly for complex or poorly defined problems. This is exactly the case in the problem at hand, finding the optimal weight values of and using the fitness function with the overall prediction error as output while also running a machine learning algorithm. Second, GAs do not need derivatives or extra information during the optimization process; they obtain the fitness score directly from the objective function [45, 46, 47]. Finally, it is straightforward to implement GAs, especially since we are also working on predicting CGM values with a CNN-LSTM network. Based on these considerations, we define our genetic algorithm’s terminologies as follows:
Genome or Gene: A real number between [1, 10] representing a weight.
Chromosome or Individual Solution: A pair of weights .
Population: A pool of chromosomes of the size of N, each representing a candidate solution.
Crossover: In each generation, crossover is performed by averaging the values of two randomly selected parents to create a new child chromosome.
Mutation: The child chromosome undergoes mutation, where a small random perturbation is added to its values. This ensures diversity in the population.
Selection: The best individuals from the current population are combined with the new offspring to form the next generation, preserving strong solutions while introducing variations.
Fitness Function: The fitness function evaluates how well each chromosome performs by calculating the total error using the custom loss function. The fitness score is simply the value of this total error; lower fitness scores indicate better solutions. In each generation, the best scores are recorded, and the individuals with the lowest fitness scores are selected as the best individuals.
By iterating through multiple generations, the genetic algorithm refines the weight parameters, aiming to find the pair that minimizes the error in critical regions. This process balances the need for generalization across all patients to optimize performance in dysglycemic regions. Algorithm 1 provides a high-level pseudo-code overview of the genetic algorithm process proposed to identify the best pair of weights. Fig. 4 illustrates the complete methodology for the design and evaluation of GLIMMER.
IV Evaluation Approach
IV-A Dataset
IV-A1 OhioT1DM Dataset
@@ IV-A2 AZT1D Dataset
In addition to using the OhioT1DM dataset, we also gathered data from 25 patients with T1D on AID systems who visited the endocrinology clinic at the Mayo Clinic in Scottsdale, AZ, between December 2023 and April 2024 for their regular appointments. Informed consents were taken from the recruited participants under the study named Machine Learning Design to Predict and Manage Postprandial Hyperglycemia in Patients with Type 1 Diabetes (IRB # 23-003065). For each patient, the data includes, on average, 26 days of recordings collected in real-world settings, featuring CGM signals captured using Dexcom G6 Pro, insulin logs, meal carbohydrate sizes, and device modes (regular/sleep/exercise) obtained using Tandem t:slim X2 insulin pump. This dataset offers comprehensive insights into diabetes management by documenting various bolus events. These include standard boluses for meal-related carbohydrate intake, correction boluses for glucose adjustments, and automatic corrections made by the pump’s algorithm. Each event is meticulously recorded, detailing insulin doses, target blood glucose levels, and user adjustments. This group of participants consists of 13 females and 12 males, aged between 27 and 80 years, with an average age of 59 years. This dataset contained 320,488 CGM entries, covering 26,707 hours of monitoring data.
IV-B Data Preparation
The OhioT1DM dataset consists of 24 files containing data from 12 patients for training, validation, and testing. For each individual, there exist two files: one for training and one for testing [48]. To prepare the data for use in our evaluation, we chronologically partition the training file, using the first 80% for training and the remaining 20% for validation. This setup ensures that the validation dataset effectively evaluates the model after hyperparameter tuning without any risk of data leakage. The entire testing file is reserved for testing purposes. We apply a similar approach to the AZT1D dataset, first splitting the data of each patient into 80% for training and 20% for testing and then dividing the training set into 80% for training and 20% for validation. This method of partitioning has been widely utilized in previous studies on blood glucose prediction [28, 49]. In addition to the CGM values, basal insulin, bolus insulin, and carbohydrate amounts are included in the datasets.
In addition to existing data that are used as features, we also added other features that might improve the model performance. Using the 200-period moving average as a feature in forecasting CGM values can be highly beneficial. In the context of CGM data, where glucose levels can fluctuate due to various factors like meals, physical activity, and stress, the 200-period moving average helps smooth out these short-term fluctuations. This smoothing effect allows the model to better capture the underlying long-term trends in glucose levels, which are crucial for making accurate predictions [50, 51]. By focusing on the long-term trend, the model is less likely to overfit short-term spikes or drops, thereby improving its generalization to new data. Another feature we implemented involves assigning a class label to CGM values. As discussed in Section III, we can categorize CGM values based on the hypoglycemia and hyperglycemia thresholds. We assign integer values (1, 2, or 3) to represent different regions: hypoglycemia, normal, and hyperglycemia. These region identifiers serve as an input feature to the model. Overall, the data preprocessing involved gathering 6 input features and applying a sliding window to the multivariate sequences, as shown in Fig. 5.
@@ IV-B Data Preparation
IV-C Architecture Configuration
We reviewed previous studies to determine common configurations for CNN-LSTM models, focusing on the number of layers and LSTM units. From the insights gained in prior studies [52, 53], we identified a range of configurations, noting that these references utilized 1 to 4 convolutional layers and LSTM units in the range of 8 to 128. This understanding guided us in designing our experiment to systematically investigate the number of layers and units for our CNN-LSTM model. We utilized a subset of the dataset, dividing it into 80% for training and 20% for validation. We tested various configurations, including 1, 2, or 3 convolutional layers with 32, 16, and 8 filters and kernel sizes of 4 and 1 or 2 LSTM layers with 4, 8, 16, and 32 units. After each convolutional or LSTM layer, we applied a dropout layer with a rate of 0.1. The results of these experiments are shown in Table I. The optimal model configuration includes three 1D convolutional layers with kernel sizes of 4 and filter counts of 32, 16, and 8, respectively, to extract features from the input data. These features are then passed to an LSTM layer with 8 units for sequential processing. A flattening layer follows, preparing the features for a dense layer with a ReLU activation function that generates the final output. The proposed architecture is illustrated in Fig. 6.
IV-D Custom Loss Function Parameters Configuration
We designed an experiment and ran Algorithm 1 to find the best weights for our custom loss function. For each patient, we created a random population of 20 candidates, each with weight pairs in the range , and evolved them over 25 generations following the methodology described in the Error Weights section. To simplify the process, we set to 1 and focused on optimizing the two parameters, and . In each generation, we evaluated the performance of each candidate by setting the weights for the custom loss function and compiling and running our CNN-LSTM model. We then calculated the RMSE on the validation dataset, which served as the fitness score for each candidate. This process involved selecting the top-performing candidates to form the basis for the next generation, applying crossover and mutation operations to generate new candidates, and repeating the evaluation process. Ultimately, we identified the best weights for each patient. However, calculating the optimal weights for each individual patient is not feasible in a practical setting due to the time-consuming nature of running the genetic algorithm for each case. Instead, to provide a practical solution, we computed the average weights across all patients, resulting in . This average provides a balanced approach that generalizes well across patients while avoiding the inefficiencies of individual optimization. Fig. 7 shows the optimal weights for each patient alongside the average pair of weights. This method ensures a reasonable approximation of the custom loss function’s performance without the impracticality of personalized optimization for each patient.
IV-E Experimental Setup
After data preparation and finalizing the GLIMMER model by setting up the parameters of the architecture and custom loss function, we conducted extensive experiments using a batch size of 48, over 30 epochs, with a prediction horizon (PH) of 60 minutes. To minimize the impact of randomness, we conducted 10 iterations of the experiments for each patient using a unique seed number for each run and reported the average results of these trials. The code used for these experiments is available for other researchers to reproduce our findings. All experiments were performed on an Apple M3 Pro chip featuring a 12-core CPU, an 18-core GPU, a 16-core Neural Engine, and 18 GB of unified memory.
IV-F Evaluation Metrics
@@
IV-F1 Root Mean Square Error (RMSE)
@@
IV-F2 Mean Absolute Error (MAE)
@@ IV-F5 F1 Score
IV-F6 Clarke Error Grid Analysis
@@
Region D: Represents predictions where critical hypoglycemia or hyperglycemia might be missed, posing a potential danger.
@@
V Results
Tables II and III summarize the performance comparison of GLIMMER with other methods for the OhioT1DM and AZT1D datasets, both using a PH of 60 minutes. We selected the most recent studies that employed a variety of classical and machine learning models, including Fast-Adaptive and Confident Neural Network (FCNN) [28], CRNN [30], Bi-LSTM [36], transformer models [55], Random Forest Regression (RFR) [56, 28], Support Vector Regression (SVR) [57], and Autoregressive Integrated Moving Average (ARIMA) [58]. Additionally, we included the GLIMMER model without modifications, represented as a basic CNN-LSTM, to highlight the effects of our enhancements, which incorporate two crafted features and a custom loss function.
@@ V Results
The results include RMSE and MAE, presented as (Mean ± Standard Deviation), along with CEG reports, which serve as standard error metrics for predictions. For both RMSE and MAE, lower values indicate better performance. As shown in Table II, GLIMMER demonstrates outstanding performance compared to other models, achieving a 23% improvement in RMSE and a 31% improvement in MAE relative to the best-reported errors. The basic CNN-LSTM results further indicate its potential as a viable candidate for analyses where other models may not be applicable. In the CEG analysis, higher values in Region A are preferable, while lower values are desirable in other regions. Notably, GLIMMER’s predictions achieve 85% within Region A, yielding a 15% improvement compared to FCNN, and it also maintains one of the lowest values in other regions.
@@
Table III presents the results for GLIMMER, and the basic CNN-LSTM model applied to the AZT1D dataset. GLIMMER again outperforms the basic CNN-LSTM model, achieving a 24% improvement in RMSE and a 28% improvement in MAE. Since this dataset has been recently collected and is not publicly available, we could only generate results for these two models, leaving evaluations for other methods to future researchers. However, given the close results in Table II for the basic CNN-LSTM and leading models like FCNN and CRNN, we anticipate similar outcomes.
To ensure a fair comparison, we did not include the MLR model in Table II, as this model employs a multi-model approach that requires training a separate model for each PH. In our case, with a PH of 60 minutes, this necessitates the creation of 12 individual models to predict the next 5, 10, 15, …, 55, and 60 minutes. In contrast, all the methods presented in Table II utilize a single model for their predictions, and the MLR model did not report CEG analysis in their studies. While they achieved an RMSE of 24.58 mg/dL and an MAE of 17.42, which are better than those of some other models, GLIMMER still outperformed them.
We visualized the results of the CEG analysis for both the OhioT1DM and AZT1D datasets in Fig. 8. The enhanced prediction accuracy of the GLIMMER model compared to the basic CNN-LSTM is evident, demonstrating its ability to accurately forecast blood glucose levels even in critical regions, which is crucial for clinical analysis and real-world applications. Additionally, Fig. 9 illustrates the glucose level predictions of GLIMMER compared to the basic CNN-LSTM. This figure highlights GLIMMER’s ability to accurately predict blood glucose levels, particularly during peaks, due to the integration of the custom loss function. The consistent results across the OhioT1DM and AZT1D datasets further demonstrate that GLIMMER is robust and generalizable, making it suitable for use with various datasets.
@@
In our analysis, we calculated error metrics separately for normal glucose levels and dysglycemic regions, providing a clearer understanding of the model’s performance across different glucose conditions. This detailed breakdown highlights the model’s strengths and identifies areas for improvement in managing varying glucose states. Precision and recall offer valuable insights: high precision indicates that the model accurately predicts dysglycemia, reducing false alarms, while high recall shows that the model effectively detects dysglycemic events, minimizing missed occurrences.
Tables IV and V present these metrics for the OhioT1DM and AZT1D datasets, comparing the performance of the basic CNN-LSTM model and GLIMMER. These metrics, along with RMSE and MAE, provide a comprehensive assessment of the model’s reliability and effectiveness in supporting diabetes management and patient safety. In the OhioT1DM dataset, GLIMMER significantly improves hypoglycemia detection, increasing recall from 16% to 42%, meaning it captures more true low blood sugar events, which is vital for patient safety. The F1 score for hypoglycemia rises from 35% to 44%, reflecting a better balance between detecting true events and minimizing false positives. For hyperglycemia, GLIMMER also improves recall by 10%, from 78% to 86%, ensuring more high blood sugar episodes are detected. With a 5% improvement in precision, it reduces false alarms, making the model more reliable and user-friendly in managing glucose levels. In the AZT1D dataset, GLIMMER shows notable gains in both recall and precision for hypoglycemia detection. Recall improves from 2% to 13%, meaning it catches more low blood sugar events, while precision rises from 34% to 47%, reducing unnecessary alerts. For hyperglycemia, GLIMMER’s recall jumps from 48% to 73%, a 52% improvement, allowing it to detect more high glucose events. The model also enhances precision by 15%, ensuring better accuracy in its predictions, making it a more effective tool for managing critical dysglycemic events.
@@
VI Discussion
@@
Our experiments demonstrated that GLIMMER achieves superior performance and accuracy as a predictive model, surpassing the state-of-the-art models. Additionally, CEG analysis validated its reliability in detecting and forecasting dysglycemic events. Recent work by Annuzzi et al. [59] explored how certain features affect blood glucose prediction using XAI methodologies. They reported an RMSE of 24.47 ± 4.27 on the AI4PG dataset, with a PH of 60 minutes. Their model depends on detailed data inputs, including preprandial blood glucose levels, insulin dosages, and meal-related factors such as energy intake, macro-nutrients, glycemic index, and glycemic load. However, we note that in real-world scenarios, obtaining these details can be challenging, as it often requires manual logging of meal and nutritional information. GLIMMER, on the other hand, achieves strong predictions using only features like CGM data, bolus and basal insulin levels, and carbohydrate amounts that are commonly obtained in automated insulin delivery systems by default. As a result, GLIMMER does not impose any additional data collection burden on the patients beyond what the standard of care requires.
@@
Despite these achievements, certain limitations remain. While GLIMMER shows improvement in F1 score, precision, and recall over the baseline CNN-LSTM model, its hypoglycemia prediction performance is still limited. One reason for this limitation is the sparsity of hypoglycemic events in both the OhioT1DM and AZT1D datasets used in this study. The scarcity of these events means the algorithm has fewer examples from which to learn hypoglycemia patterns effectively. Additionally, we were unable to compare GLIMMER to other models on the AZT1D dataset because the source codes for these methods are not publicly available. In future work, we plan to implement and evaluate additional models to achieve a more comprehensive assessment.
Although the core architecture of GLIMMER is CNN-LSTM, the methodologies presented in this article, including the proposed custom loss function and the optimization algorithm, are model-agnostic and can be adapted to other neural network models as well. Future work will focus on designing more advanced architectures, such as attention-based models, to explore GLIMMER’s potential within more complex time-series forecasting frameworks. We also aim to extend the prediction horizon and incorporate long-term features, which could improve accuracy over longer time periods. As shown in Fig. 1, integrating GLIMMER with automated insulin delivery devices and developing a smartphone application to alert patients and physicians about potential dysglycemic events could have transformative effects. This application would provide projections for the next hour of blood glucose levels, acting as a preventive tool and allowing for a comparison of GLIMMER’s effectiveness against current methods.
VII Conclusion

### v2

@@ 3.2 Custom Loss Function
3.3 Error Weights
We classify CGM values into regions based on the thresholds discussed earlier. Next, we finalize the custom loss function by determining the optimal weights for each region, focusing on dysglycemia. To simplify optimization, we fix , reducing the task to finding and . To find these weights, we use a Genetic Algorithm (GA) [39], a robust optimization method well-suited for problems where the fitness function involves training a model and evaluating its error. GAs require no derivatives or assumptions about the search space [39, 40, 41], and their flexibility makes them ideal for this task. In our implementation, the GA optimizes and for each patient by evolving a population of candidate weight pairs. This ensures the weights are adapted to the clinical relevance of each region and tailored to each patient. The procedure is as follows:
1.
For each patient, initialize a random population of weight pairs , sampled uniformly from the range .
2.
Repeat the following for generations:
Evaluate the fitness of each individual by training the model with its weights and computing the total prediction error on validation data.
Select the top individuals with the lowest errors as parents.
Generate offspring as follows:
i.
Randomly select two parents and average their weights to create a child.
ii.
Add Gaussian noise () to the child weights to introduce variation.
iii.
Clip the child weights to ensure they remain within .
i.
Form the next generation by combining the parents and offspring.
3.
After generations, select the weight pair with the lowest validation error as the optimal .
In summary, our methodology combines three key components to enhance prediction performance: selecting a suitable model architecture, designing a custom loss function that emphasizes clinically critical regions, and using a genetic algorithm to tune the loss function’s weights. Figure 3 illustrates the complete methodology for the design and evaluation of GLIMMER. In the following sections, we describe the feature engineering that contributes to improved model performance, as well as the experiments conducted to determine the optimal CNN-LSTM architecture and custom loss function weights.
4 Data
4.1 OhioT1DM Dataset
@@ 4.2 AZT1D Dataset
Alongside the OhioT1DM dataset, we introduced and published a new dataset named AZT1D [17]. We collected data from 25 patients with T1D using AID systems who visited the endocrinology clinic at the Mayo Clinic in Scottsdale, AZ, between December 2023 and April 2024 for their regular appointments. The participants included 13 females and 12 males, aged between 27 and 80 years, with an average age of 59 years. Informed consent was obtained from all participants under the study protocol (IRB #23-003065). For each patient, the dataset contains an average of 26 days of real-world recordings, including CGM signals captured with Dexcom G6 Pro, insulin logs, meal carbohydrate sizes, and device modes (regular/sleep/exercise) recorded from the Tandem t:slim X2 insulin pump. In total, the dataset comprises 320,488 CGM entries spanning approximately 26,707 hours of monitoring data.
4.3 Data Preparation
The OhioT1DM dataset consists of 24 files from 12 patients, with two files per patient: one for training and one for testing [42]. To avoid overfitting and to enable hyperparameter tuning, we further split the training file chronologically, using the first 80% for training and the remaining 20% for validation. The separate testing file is kept intact and used exclusively for final evaluation.
For the AZT1D dataset [17], which does not come pre‑partitioned, we applied a similar strategy: each patient’s data was split into 80% for training and 20% for testing, and then the training portion was again split into 80% training and 20% validation.
In addition to CGM values, basal insulin, bolus insulin, and carbohydrate amounts, we crafted two additional features to enhance model performance. First, we computed the 200‑period moving average of CGM values to smooth short‑term fluctuations from meals, activity, or stress, allowing the model to focus on the long‑term trend relevant for predictions [43, 44]. This trend emphasis reduces overfitting to transient spikes or drops. Second, we assigned each CGM value a class label (hypoglycemia, normal, hyperglycemia) based on thresholds, as detailed in Section 3, helping the model recognize clinically important ranges. In total, six input features were used in a sliding‑window format (Figure 4).
5 Experimental Setup
5.1 Architecture Configuration
We reviewed previous studies to identify common configurations for CNN‑LSTM models, which typically use 1–4 convolutional layers and 8–128 LSTM units [30, 45]. Guided by these insights, we designed an experiment to systematically evaluate different layer and unit configurations. We used a subset of the dataset, split into 80% for training and 20% for validation. The results of these experiments are summarized in Table 2. The final architecture, shown in Figure 5, consists of three 1D convolutional layers (kernel size 4; 32, 16, and 8 filters), an LSTM layer with 8 units, a flattening layer, and a dense layer with ReLU activation.
@@ 5.1 Architecture Configuration
5.2 Custom Loss Function Parameters Configuration
We conducted an experiment using the genetic algorithm described in Section 3 to determine the optimal weights for our custom loss function. In each generation, the performance of each candidate was evaluated by applying its weights to the custom loss function, compiling, and running the CNN‑LSTM model. The RMSE on the validation dataset was then computed and used as the fitness score. Through this process, we identified the best weights for each patient. However, calculating personalized weights for every patient is impractical in real‑world settings due to the computational cost of running the genetic algorithm repeatedly. To address this, we computed the average weights across all patients, yielding . This averaged configuration offers a balanced solution that generalizes well while avoiding the inefficiencies of per‑patient optimization, providing a practical approximation of the custom loss function’s performance.
5.3 Evaluation Metrics
@@
5.3.1 Root Mean Square Error (RMSE)
@@
5.3.2 Mean Absolute Error (MAE)
@@ 5.3.5 F1 Score
5.3.6 Clarke Error Grid Analysis
@@
4.
Region D: Represents predictions where critical hypoglycemia or hyperglycemia might be missed.
5.
@@
6 Results
After data preparation and finalizing the GLIMMER model by setting up the parameters of the architecture and custom loss function, we conducted extensive experiments using a batch size of 48, over 30 epochs, with a prediction horizon (PH) of 60 minutes. To minimize the impact of randomness, we conducted 10 iterations of the experiments for each patient using a unique seed number for each run and reported the average results of these trials. All experiments were performed on an Apple M3 Pro chip featuring a 12-core CPU, an 18-core GPU, a 16-core Neural Engine, and 18 GB of unified memory.
6.1 Overall Performance
Tables 3 and 4 summarize the performance comparison of GLIMMER with other methods for the OhioT1DM and AZT1D datasets, both using a PH of 60 minutes. We selected the most recent studies that employed a variety of classical and machine learning models, including Fast-Adaptive and Confident Neural Network (FCNN) [16], CRNN [21], Bi-LSTM [27], transformer models [32], Random Forest Regression (RFR) [34, 16], Support Vector Regression (SVR) [47], and Autoregressive Integrated Moving Average (ARIMA) [33]. Additionally, we included the original GLIMMER model, implemented as a basic CNN‑LSTM, to highlight the impact of our enhancements. These enhancements incorporate CGM region classes and moving averages as additional features, a custom loss function optimized with a genetic algorithm, improved architecture design, and tuned hyperparameters.
The results include RMSE and MAE, reported as (Mean ± Standard Deviation), along with CEG metrics to assess prediction errors. Lower RMSE and MAE indicate better performance. As shown in Table 3, GLIMMER achieves a 23% improvement in RMSE and 31% in MAE compared to the best‑reported errors. The basic CNN‑LSTM also performs well, making it a viable option when other models are unsuitable. In the CEG analysis, higher values in Region A and lower in other regions are desirable. GLIMMER attains 85% in Region A, a 15% improvement over FCNN, while maintaining low values elsewhere.
@@ 6.1 Overall Performance
Table 4 shows results for GLIMMER and the basic CNN‑LSTM on the AZT1D dataset. Again, GLIMMER outperforms the basic CNN‑LSTM, with a 24% improvement in RMSE and 28% in MAE. Since this dataset is newly collected and not publicly available, we report results only for these two models, leaving broader comparisons to future work. Given the close results on the OhioT1DM dataset between CNN‑LSTM and other models like FCNN and CRNN, similar trends are expected here.
To ensure a fair comparison, we excluded the MLR model from Table 3, as it requires training a separate model for each prediction horizon (PH). At a PH of 60 minutes, this entails 12 individual models for predicting the next 5, 10, …, up to 60 minutes. In contrast, the models in Table 3 use a single model for all horizons. Moreover, the MLR study did not report CEG results. While its RMSE (24.58 mg/dL) and MAE (17.42) surpass some models, GLIMMER still outperforms it.
Figure 6 visualizes the CEG analysis results for both the OhioT1DM and AZT1D datasets, complementing the numeric results reported in Tables 3 and 4. The plots compare the basic CNN‑LSTM and the GLIMMER model across datasets, showing how predictions are distributed across the CEG regions. GLIMMER’s predictions are more concentrated in Region A and closer to the diagonal line, indicating better clinical accuracy and fewer large errors, particularly in critical regions. This visualization highlights GLIMMER’s improved ability to forecast blood glucose levels reliably, which is crucial for both clinical evaluation and practical deployment.
6.2 Regional and Event-Specific Performance
@@
In our analysis, we calculated error metrics separately for normal glucose levels and dysglycemic regions, providing a clearer understanding of the model’s performance across different glucose conditions. This detailed breakdown highlights the model’s strengths and identifies areas for improvement in managing varying glucose states. Precision and recall offer valuable insights: high precision indicates that the model accurately predicts dysglycemia, reducing false alarms, while high recall shows that the model effectively detects dysglycemic events, minimizing missed occurrences. Tables 5 and 6 present these metrics for the OhioT1DM and AZT1D datasets, comparing the performance of the basic CNN-LSTM model and GLIMMER. These metrics, along with RMSE and MAE, provide a comprehensive assessment of the model’s reliability and effectiveness in supporting diabetes management and patient safety.
@@
In the OhioT1DM dataset, GLIMMER significantly improves hypoglycemia detection, increasing recall from 16% to 42%, meaning it captures more true low blood sugar events, which is vital for patient safety. The F1 score for hypoglycemia rises from 35% to 44%, reflecting a better balance between detecting true events and minimizing false positives. For hyperglycemia, GLIMMER also improves recall by 10%, from 78% to 86%, ensuring more high blood sugar episodes are detected. With a 5% improvement in precision, it reduces false alarms, making the model more reliable and user-friendly in managing glucose levels.
In the AZT1D dataset, GLIMMER shows notable gains in both recall and precision for hypoglycemia detection. Recall improves from 2% to 13%, meaning it catches more low blood sugar events, while precision rises from 34% to 47%, reducing unnecessary alerts. For hyperglycemia, GLIMMER’s recall jumps from 48% to 73%, a 52% improvement, allowing it to detect more high glucose events. The model also enhances precision by 15%, ensuring better accuracy in its predictions, making it a more effective tool for managing critical dysglycemic events.
To better interpret the improvements presented in Tables 5 and 6, Figure 7 highlights the advantages of GLIMMER over the baseline CNN-LSTM model, particularly in clinically significant regions. The figure demonstrates that GLIMMER more accurately captures blood glucose fluctuations—especially during post-meal spikes—due to the integration of a targeted loss function designed to emphasize errors in dysglycemic ranges. These improvements are clinically meaningful, as they may allow for earlier detection of postprandial hyperglycemia and more timely corrective action. Such capability addresses a critical limitation in current AID systems, which often react too late or deliver insufficient boluses after meals. The model’s consistent performance across both the OhioT1DM and AZT1D datasets further supports its robustness and generalizability.
7 Discussion
@@
Our experiments demonstrated that GLIMMER achieves superior performance and accuracy as a predictive model, surpassing the state-of-the-art models. Additionally, CEG analysis validated its reliability in detecting and forecasting dysglycemic events. Recent work by Annuzzi et al. [48] explored how certain features affect blood glucose prediction using XAI (Explainable AI) methodologies. They reported an RMSE of 24.47 4.27 on the AI4PG dataset, with a prediction horizon of 60 minutes. Their model depends on detailed data inputs, including preprandial blood glucose levels, insulin dosages, and meal-related factors such as energy intake, macro-nutrients, glycemic index, and glycemic load. However, we note that in real-world scenarios, obtaining these details can be challenging, as it often requires manual logging of meal and nutritional information. GLIMMER, on the other hand, achieves strong predictions using only features like CGM data, bolus and basal insulin levels, and carbohydrate amounts that are commonly obtained in automated insulin delivery systems by default. As a result, GLIMMER does not impose any additional data collection burden on the patients beyond what the standard of care requires.
Although similar works, such as that of Cichosz et al. (2021) [49], introduced the concept of error-weighted glucose prediction, GLIMMER advances this approach in several important ways. Rather than relying on fixed penalty weights, our method employs a genetic algorithm to learn data-driven weights for a custom loss function, enabling more adaptive error prioritization. GLIMMER also places explicit emphasis on minimizing prediction errors in clinically significant dysglycemic regions—hypoglycemia and hyperglycemia—where accurate forecasts are most critical. In terms of architecture, we leverage a CNN-LSTM model tailored to capture the temporal dynamics and multivariate structure of glucose data, offering greater representational power compared to shallow feed-forward networks used in earlier work. This is further enhanced by domain-specific feature engineering, designed to reflect physiological patterns underlying glucose fluctuations.
@@
In light of the contributions presented, some limitations still warrant discussion. While GLIMMER shows improvement in F1 score, precision, and recall over the baseline CNN-LSTM model, its hypoglycemia prediction performance is still limited. One reason for this limitation is the sparsity of hypoglycemic events in both the OhioT1DM and AZT1D datasets used in this study. The scarcity of these events means the algorithm has fewer examples from which to learn hypoglycemia patterns effectively. Additionally, we were unable to compare GLIMMER to other models on the AZT1D dataset because the source codes for these methods are not publicly available. In future work, we plan to implement and evaluate additional models to achieve a more comprehensive assessment.
The performance of glucose prediction models like GLIMMER is inherently dependent on the accuracy and reliability of glucose sensing technologies. Currently, CGMs represent the clinical standard, providing near real-time interstitial glucose readings with steadily improving accuracy and usability [50]. CGMs remain central to most automated insulin delivery systems and were the primary input source in our study. However, they are still constrained by physiological lag time, sensor calibration needs, and signal noise. Emerging biosensing methods—such as sweat-based detection [51], breath analysis [52], reverse iontophoresis [53], and optical spectroscopy [54]—offer potential for non-invasive or faster sensing, but most remain experimental and lack the clinical validation required for integration into routine care. Improvements in biosensor technology could reduce latency, improve measurement fidelity, and ultimately enhance the quality and reliability of glucose forecasts. As these innovations mature, they may open new opportunities to refine models like GLIMMER or extend them to entirely new care settings.
Despite GLIMMER’s reliance on a CNN-LSTM structure, the underlying methodologies introduced, such as the custom loss function and optimization algorithm, are generalizable and can be applied to other architectures. As part of future work, we plan to design more advanced forecasting models, such as attention-based or transformer-based networks, to explore GLIMMER’s performance within more complex time-series frameworks. We also aim to extend the prediction horizon up to 120 minutes and incorporate longer-term features to enhance accuracy over extended periods. As illustrated in Figure 1, integrating GLIMMER with automated insulin delivery systems and deploying a companion smartphone application could enable real-time alerts for patients and clinicians, providing two-hour-ahead blood glucose projections and supporting early intervention. Additionally, we plan to expand the AZT1D dataset from 25 to 100 patients, further strengthening the generalizability of our findings and enabling richer model development and evaluation.
8 Conclusion